OpenStudy (anonymous):

Find the length of the base of a square pyramid if the volume is 1,024 cubic inches and has a height of 12 inches. Use 3.14 for pi and round your answer to the nearest hundredth. 4 inches 8 inches 16 inches 24 inches

OpenStudy (anonymous):

@ganeshie8 @phi @jojo4eva Please help

OpenStudy (phi):

what is the formula for the volume of a pyramid?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

that looks like a cone. You want the formula for a square pyramid.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

OK, can you post the equation for the volume of a square pyramid?

OpenStudy (anonymous):

1/3 times b squared times H

OpenStudy (phi):

you show write down the entire equation, because we are going to use it to answer the question.

OpenStudy (phi):

and use math not words

OpenStudy (anonymous):

(1/3) b h

OpenStudy (phi):

an equation has an = sign in it

OpenStudy (anonymous):

V=(1/3) b h

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Now we are ready. they tell you volume is 1,024 height of 12 inches replace the letters in your equation with those numbers. what do you get ?

OpenStudy (phi):

and somehow your b^2 lost its 2

OpenStudy (anonymous):

1,024=(1/3) B2 12

OpenStudy (phi):

you should write B^2 (btw B is the side of the square... what you are trying to find) I would have used x or s for the side of the square... your equation is \[ 1024 = \frac{1}{3} b^2 \cdot 12 \] now you have to solve for b. First, multiply both sides by 3. what do you get ?

OpenStudy (anonymous):

12 times 3 : 36

OpenStudy (phi):

No, we don't want to do that. How about this: what is 12/3 ?

OpenStudy (anonymous):

4

OpenStudy (phi):

can you re-write your equation, using that simplification that 12/3 is 4: \[ 1024 = \frac{1}{3} b^2 \cdot 12 \]

OpenStudy (anonymous):

1024=B2*4

OpenStudy (phi):

ok, but get used to writing B^2 (because B2 really means B times 2) now divide both sides by 4.

OpenStudy (phi):

like this \[ \frac{1024}{4} = \frac{4b^2}{4} \] simplify both sides. on the right side 4/4 is 1 on the left side 1024/4 is 256 to find b (the answer) take the square root of both sides. (remember that the sqr(b^2) is b )

OpenStudy (anonymous):

so my answer would be 16

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you can you help me with something else please